OpenStudy (anonymous):

An investor invested a total of $2,300 in two mutual funds. One fund earned a 9% profit while the other earned a 5% profit. If the investor's total profit was $159, how much was invested in each mutual fund?

OpenStudy (anonymous):

The total profit is 163. Not 159. I'm not sure how this works, I have been stuck on this problem for about an hour

OpenStudy (anonymous):

just wanna make sure is 2300 invested in each one? this is what i have so far x = amt invested in 1st fund with 9% y = amt invested in 2nd x + y = 2300, x(1.09)= some amount made from first fund(a) y(1.05) = some amount made from 2nd(b) the total profit is 163 total cash is 163

OpenStudy (anonymous):

sorry nvm i try to figure it out

OpenStudy (anonymous):

2300 is the total investment. 9% and 5% was invested in the 2300. The total profit of each investment was 163. How much was invested at 9% and how much was invested at 5%. I got 1400 at 9% and 900 at 5%?

OpenStudy (anonymous):

but that is 171. I keep getting stuck

OpenStudy (anonymous):

lets try this x(0.09) + y(0.05) = 163 x = (163 - 0.05y) / 0.09 (163 - 0.05y) / 0.09 + y = 2300 ((163-0.05y) + 0.09y)/0.09 = 2300 163 - 0.05y + 0.09y = 207 163 + 0.04y = 207 0.04y = 44 y = 1100 x + 1100 = 2300 x = 1200

OpenStudy (anonymous):

x = amount invested in the first one y = amount invested in 2nd one to check x + y = 2300 1200 + 1100 = 2300 total invested 1200(9%) = 108 1100(5%) = 55 total profit = 108 + 55 = 163

OpenStudy (anonymous):

That is the formula I have been looking for! Thank you! Okay I remember how to finish the problem.

OpenStudy (anonymous):

np glad to help